# BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v2

@@ Abstract
6D pose estimation aims at determining the pose of the object that best explains the camera observation. The unique solution for a non-symmetrical object can turn into a multi-modal pose distribution for a symmetrical object or when occlusions of symmetry-breaking elements happen, depending on the viewpoint. Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries, whereas they should be defined per-image to account for the camera viewpoint. We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the visibility of the object surface in the image to correctly determine the visual ambiguities. Second, given this improved ground truth, we re-evaluate the state-of-the-art single pose methods and show that this greatly modifies the ranking of these methods. Third, as some recent works focus on estimating the complete set of solutions, we derive a precision/recall formulation to evaluate them against our image-wise distribution ground truth, making it the first benchmark for pose distribution methods on real images. We will make our annotations for the T-LESS dataset and our code publicly available.
@@ 1 Introduction
In this paper, we first propose a method that automatically annotates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions. Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models. While we demonstrate it on T-LESS [13], our method is fully automatic and can be applied to other datasets as demonstrated on YCB-V [47].
@@
    A comprehensive re-evaluation of 6-DOF single pose estimation methods related to the re-annotation of the T-LESS [13] ground-truth;
@@ 2 Related Work
Object 6D pose estimation aims at determining an object pose that best explains the camera observation. While the solution is unique when the object is non-symmetrical and not occluded, multiple equivalent possible solutions arise when the object is symmetrical or when the occlusions prevent the observation of symmetry-breaking elements.
While the state of the art mostly focused on estimating a single pose [6, 38, 15, 23, 41, 44, 8, 40], some recent works focus on estimating the complete set of solutions [29, 36, 33, 22, 30, 18, 9, 19]. As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that return a single pose per image.
@@
Various techniques have been introduced to annotate image datasets with their 6D poses. For synthetic datasets [2, 43], the ground truth pose is directly available, while, for real datasets, this pose is usually determined with the help of user interaction [7, 14, 5], markers [10, 11, 1], or a robotic arm [45, 31], and refined with ICP in the case of an RGB-D camera [7, 14, 45, 31].
@@
For non-symmetrical objects, recovering the complete set of solutions is much more challenging since symmetries may arise from the non-visibility of disambiguating parts of the object, those non-visibilities being induced by the viewpoint (self-occlusion), by other elements of the scene (external occlusion). The transformation set that relates the different solutions is then specific to each image and cannot be pre-computed from the 3D object model.
To our knowledge, no method has been proposed to determine this ground-truth transformation set per image. Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set. The latter is usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g. surface deviation tolerance of 1.5cm or 10% of the object diameter [16]).
@@
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [12]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 47, 43]. In practice, the most commonly used accuracy measures are :
@@
    Average Distance metric (ADD), and variations for symmetrical objects (ADD-S [47], ADD-H [43]), target to measure mean Euclidean error between surface points of the object transformed with the estimated pose and the ground truth pose, but are replaced by MSSD now.
@@
    Maximum Symmetry-Aware Surface Distance (MSSD) [16] which is similar to ADD but considers the maximal error instead of the mean error and symmetry management. MSSD provides a 3D error, useful to robotics applications.
@@
    Maximum Symmetry-Aware Projection Distance (MSPD) [16] measures the maximum reprojection error between the projection of the visible points with the ground-truth pose and their corresponding projection with the pose to evaluate and symmetry management. MSPD provides a reprojection error, useful to augmented reality applications.
@@ 3 Method
As shown in Figure 2, the probability of a pose explaining a given image of the object is related to the visibility of the different points on the surface of the object.
More formally, given an image $I$ and a ground truth pose $P_{\text{GT}}$ for a target object in this image, the objective of our method is to recover the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that transform the ground truth pose $P_{\text{GT}}$ into another pose that results in the same appearance of the object. In the following, we will call set ${\cal T}$ the symmetry pattern of image $I$.
Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non-ambiguous pose or can include multiple transformations in addition to the identity. As introduced in [12], ${\cal T}$ can be defined as partial $\epsilon$-symmetries of ${\cal V}(M)$:
@@
where $d$ measures the mean deviation between the points of ${\cal V}(M)$ and $M$, $\epsilon$ is the maximal surface deviation tolerance and * represents the application of the transformation $T_{i}$ onto the point cloud ${\cal V}(M)$.
If the set ${\cal V}(M)$ of visible points is available, the symmetry pattern could in principle be approximated by sampling the rigid transformation space SE(3) and assessing whether the sample belongs to the $\epsilon$-symmetries. However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used to our knowledge.
In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured objects (Section 3.3).
@@ 3.1 Speeding up the computation of $\epsilon$-sym
Computing $\epsilon$-symmetries for each instance of image of a dataset can quickly become time consuming since it implies to sample the 6D pose space around the ground truth pose. To accelerate the computation of the $\epsilon$-symmetries, we reformulate them to introduce image-independent terms that can be pre-computed.
@@
where $v_{j}$ is a 3D point and $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
With such expression of the $\epsilon$-symmetries, it appears that elementary symmetry patterns can be pre-computed for the whole 3D object model $M$ since they are independent of the image. Only the set of visible points and the intersection of the corresponding pre-computed elementary symmetries patterns need to be computed per image.
@@ 3.2 Robust symmetries patterns intersection
A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetry pattern. Our soft intersection tolerates a threshold and keeps all the $T_{i}$ such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking elements.
@@ 3.3 Generalization to textured 3D objects
where $d_{\text{color}}$ is the distance in color-space and $\zeta$ is the color deviation tolerance.
@@ 4 Evaluating 6D Pose Distribution Predictions
Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for single pose estimation, pose error for 6D pose distribution). With our annotation method that provides the full set of 6D pose solutions even on real data, it becomes possible to evaluate both tasks on the same dataset. However, it would be beneficial if both tasks can be evaluated with comparable metrics.
We therefore propose to keep the evaluation process of single pose method unchanged, but on a more accurate ground truth, and to extend it to pose distributions evaluation. First, whereas pose error is commonly used in 6D pose distribution evaluation process [33, 20, 9], we propose to use the registration error. The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the applications. It is also the type of error used by the current gold standard benchmark for single pose [42]. Typically, we suggest to use the Maximal Surface Distance (MSD) and Maximum Projective Distance (MPD) metrics. Second, we propose to keep the metrics of Precision and Recall, but we adapt them to measure Precision over the whole estimated distribution of poses instead of a unique pose, and Recall over the whole set of solutions instead of considering a pose to be found if one of its multiple solutions was found.
@@
$\mathbf{P_{d}}(\text{Est},\text{GT})=\\ \sum\limits_{P_{\text{Est}}^{j}\in Est}p(P_{\text{Est}}^{j})\left(\min_{P_{\text{GT}}\in\text{GT}}d(P_{\text{Est}}^{j},P_{\text{GT}})<\tau_{\mathbf{d}}\right)\>,$ (6)
where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method outputs solutions without assigning a probability to each pose, such as [19], a uniform probability over the solutions is used., $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
@@
$\mathbf{R_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{GT}}^{i}\in\text{GT}}\max\left(p(\hat{P}_{\text{Est}}^{i}),\frac{1}{\text{Card}(\text{GT})}\right)[d(\hat{P}_{\text{Est}}^{i},P_{\text{GT}}^{i})<\tau_{\mathbf{d}}],\\ \text{with }\hat{P}_{\text{Est}}^{i}=\operatorname*{arg\,min}_{P_{\text{Est}}\in\text{Est}}d(P_{\text{Est}},P_{\text{GT}}^{i})\>.$ (7)
@@ 5 Experiments
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and Homebrewed [21] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths. HOPE [43] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2). The second evaluates the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [17] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3). The last experiment evaluates for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
@@ 5.1 Implementation details
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2.
@@ 5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [17] public submissions have become publicly available to allow in-depth analysis. In this experiment, we use them to compare the performance evaluations of state-of-the-art methods on T-LESS for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the mode with highest confidence. We consider their full predicted distributions in the next subsection. Since these methods are not part of the BOP challenge, we retrained them on the T-LESS train set.
@@
This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website. Following the BOP Challenge, the evaluation is based exclusively on recall. Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies.

### v3

@@ Abstract
6D pose estimation aims at determining the object pose that best explains the camera observation. The unique solution for non-ambiguous objects can turn into a multi-modal pose distribution for symmetrical objects or when occlusions of symmetry-breaking elements happen, depending on the viewpoint. Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries, whereas they should be defined per-image to account for the camera viewpoint. We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the object surface visibility in the image to correctly determine the visual ambiguities. Second, given this improved ground truth, we re-evaluate the state-of-the-art single pose methods and show that this greatly modifies the ranking of these methods. Third, as some recent works focus on estimating the complete set of solutions, we derive a precision/recall formulation to evaluate them against our image-wise distribution ground truth, making it the first benchmark for pose distribution methods on real images.
@@ 1 Introduction
In this paper, we first propose a method that automatically annotates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions. Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models. Its genericity is demonstrated on T-LESS [13] and YCB-V [48].
@@
    A comprehensive re-evaluation of 6-DOF single pose estimation methods related to the re-annotation of the T-LESS [13] and YCB-V [48] ground truths;
@@ 2 Related Work
Object 6D pose estimation aims at determining an object pose that best explains the camera observation. Multiple equivalent solutions arise when the object is symmetrical or when occlusions (external or self) prevent the observation of symmetry-breaking elements (e.g. for a cup [12, 30]).
While the state of the art mostly focused on estimating a single pose [6, 39, 15, 24, 42, 45, 8, 41], some recent works focus on estimating the complete set of solutions [30, 37, 34, 23, 31, 18, 9, 19]. As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are then required, even for methods that return a single pose per image.
@@
Various techniques have been introduced to annotate image datasets with their 6D poses. For synthetic datasets [2, 44], the ground truth pose is directly available, while, for real datasets, this pose is usually determined with the help of user interaction [7, 14, 5], markers [10, 11, 1], or a robotic arm [46, 32, 21], and refined with ICP in the case of an RGB-D camera [7, 14, 46, 32].
@@
For non-symmetrical objects, recovering the complete set of solutions is much more challenging since symmetries may arise from the non-visibility of disambiguating parts of the object, those non-visibilities being induced by the viewpoint (self-occlusion), by other elements of the scene (external occlusion), or by the lack of image resolution (resolution occlusion). The transformation set that relates the different solutions is then specific to each image and cannot be pre-computed from the 3D object model.
To our knowledge, no method has been proposed to determine this ground-truth transformation set per image. Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set, usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g. surface deviation tolerance of 1.5cm or 10% of the object diameter [16]), and cannot be used to evaluate properly multi-modal pose distribution estimates like [9, 19].
@@
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [12]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground-truths due to a symmetrically tagged object, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 48, 44]. In practice, the most commonly used accuracy measures are :
@@
    Average Distance metric (ADD) [11], and variations for symmetrical objects (ADD-S [48], ADD-H [44]), measure mean Euclidean error between estimated and ground truth surface points, but are replaced by MSSD.
@@
    Maximum Symmetry-Aware Surface Distance (MSSD) [16] , similar to ADD, considers the maximal Euclidean error with symmetry management, and provides a 3D error, useful to robotics applications.
@@
    Maximum Symmetry-Aware Projection Distance (MSPD) [16] measures the maximum reprojection error between projections of the ground truth and estimated surface points, with symmetry management, and provides an error useful for augmented reality.
@@ 3 Method
Consider a 3D object model $M$ observed from a pose $P_{\text{GT}}$. We call symmetry-pattern the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that, once combined with $P_{\text{GT}}$, generate representations that are indistinguishable from the one generated by $P_{\text{GT}}$. Depending on the object geometry $M$ and the representation type, the set ${\cal T}$ can be limited to the identity transform or can include multiple transformations in addition to the identity.
Previous work well studied the representation corresponding to a 3D surface, or a subset ${\cal V}(M)\in M$ of the 3D object surface. In such case, the symmetries pattern ${\cal T}({\cal V}(M))$ can be defined through $\epsilon$-symmetries  [12]:
@@
where $\epsilon$ is the maximal deviation tolerance, * represents the application of the transformation $T_{i}$ onto the point set ${\cal V}(M)$ and $d$ measures the max deviation between the sets:
$d(X,M)=\max_{x\in X}||x-\operatorname*{arg\,min}_{m\in M}(||x-m||_{2})||_{2}$.
In this paper, we propose to consider the case of representation corresponding to an image $I$. The symmetries pattern ${\cal T}(I)$ associated to the image $I$ corresponds to a subset of the symmetries pattern ${\cal T}({\cal V}(M))$ associated to the visible part of the object ${\cal V}(M)$ in the image. Indeed, since $\epsilon$-symmetries work in 3D space, they do not take into account the potential self or external occlusions that can be induced by the projection process that produces the image.
Consequently, to estimate ${\cal T}(I)$, our solution compute first ${\cal T}({\cal V}(M))$ then prune the pose visually-inconsistent (section 3.3). As illustrated in Figure 2, to achieve this estimation efficiently, our solution decompose the estimation of ${\cal T}({\cal V}(M))$ in a two step process: a first step of pre-computation (section 3.1) independent of ${\cal V}(M)$ and executed only once, and a second step that compute ${\cal T}({\cal V}(M))$ (section 3.2).
@@ 3.1 Speeding up the computation of $\epsilon$-sym
Computing ${\cal T}({\cal V}(M))$ for each instance in each image of a dataset can quickly become time consuming since it implies to sample the 6D pose space around the ground truth pose. This can explain why such annotation process was never used to our knowledge. We accelerate the $\epsilon$-symmetries computation by introducing terms independent from ${\cal V}(M)$ that can be pre-computed.
@@
where $v_{j}$ is a 3D point and $\epsilon\text{-sym}(v_{j},M)$ is the symmetries pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetries pattern of $v_{j}$.
With such expression of the $\epsilon$-symmetries, it appears that elementary symmetries patterns can be pre-computed for the whole 3D object model $M$ since they are independent of the image. Only the set of visible points and the intersection of the corresponding pre-computed elementary symmetries patterns need to be computed per image.
@@ 3.2 Robust symmetries patterns intersection
A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetries pattern. Our soft intersection tolerates a threshold and keeps all the $T_{i}$ such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking elements.
3.3 Refining $\epsilon$-sym: visual occlusions consideration
[…]
$Prune\big{(}D_{GT},D_{T_{i}},V_{GT},V_{T_{i}},\delta\big{)}=\\ \sum_{p\in V_{GT}\cup V_{T_{i}}}\begin{cases}0&\text{if}\;p\in V_{GT}\cap V_{T_{i}}\,\wedge\,\big{|}D_{GT}(p)-D_{T_{i}}(p)\big{|}<\delta\\ 1&\text{otherwise,}\end{cases}$ (5)
where $D_{GT}$ and $V_{GT}$ are distance image and visible mask of the ground truth, $D_{T_{i}}$ and $V_{T_{i}}$ are their $T_{i}$ counterparts, with visible masks computed from sensor depth.
@@ 3.4 Generalization to textured 3D objects
where $\zeta$ is the color deviation tolerance and with:
$d_{\text{color}}(X,M)=\max_{x\in X}d_{\text{col}}(x,\operatorname*{arg\,min}_{m\in M}(||x-m||_{2}))$,
with $d_{\text{col}}(x,m)$ the color distance between vertices $x$ and $m$.
@@ 4 Evaluating 6D Pose Distribution Predictions
Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for the first, pose error for the others). With our annotation method that provides the full set of 6D pose solutions even on real data, it becomes possible to evaluate both tasks on the same dataset. However, it would be beneficial if both tasks can be evaluated with comparable metrics.
We therefore propose to keep the evaluation process of single pose method unchanged, but on a more accurate ground truth, and to extend it to pose distributions evaluation. First, whereas pose error is commonly used in 6D pose distribution evaluation process [34, 20, 9], we propose to use the registration error. The latter takes the object 3D shape into account, making the accuracy more meaningful for most applications. It is also the type of error used by the current gold standard benchmark for single pose [43]. Typically, we suggest to use the Maximal Surface Distance (MSD) and Maximum Projective Distance (MPD) metrics. Second, we propose to keep the metrics of Precision and Recall, but we adapt them to measure Precision over the whole estimated distribution of poses instead of a unique pose, and Recall over the whole set of solutions instead of considering a pose to be found if one of its multiple solutions was found.
@@
$\mathbf{P_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{Est}}^{i}\in\text{Est}}p(P_{\text{Est}}^{i})\left(\min_{P_{\text{GT}}\in\text{GT}}\mathbf{d}(P_{\text{Est}}^{i},P_{\text{GT}})<\tau_{\mathbf{d}}\right)\>,$ (7)
where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(P_{\text{Est}}^{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method outputs solutions without assigning a probability to each pose, such as [19], a uniform probability over the solutions is used., $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
@@
$\mathbf{R_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{GT}}^{i}\in\text{GT}}\min\left(p(\hat{P}_{\text{Est}}^{i}),\frac{1}{\text{Card}(\text{GT})}\right)[\mathbf{d}(\hat{P}_{\text{Est}}^{i},P_{\text{GT}}^{i})<\tau_{\mathbf{d}}],\\ \text{with }\hat{P}_{\text{Est}}^{i}=\operatorname*{arg\,min}_{P_{\text{Est}}\in\text{Est}}\mathbf{d}(P_{\text{Est}},P_{\text{GT}}^{i})\>.$ (8)
@@ 5 Experiments
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and HB [22] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths. HOPE [44] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2). The second evaluates the impact of our re-annotation of T-LESS and YCB-V datasets on the performance evaluation of the methods of the BOP benchmark [17] on single pose estimation  (Section 5.3). The last experiment evaluates for the first time 6D pose distribution methods on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
@@ 5.1 Implementation details
Surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
@@ 5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [17] submissions have become publicly available to allow in-depth analysis. This experiment reprocess these results to compare the performance evaluations of state-of-the-art methods for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the highest confidence mode. Since these methods are not part of the BOP challenge, we retrained them.
@@
This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website. Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies.